A chemist needs 18 L of a 10% alcohol solution but only has a 30% alcohol solution. How many liters each of the 30 solution and water should be mixed to make the desired 18 liters of a 10% solution?
Answer by stanbon(75887)

A chemist needs 18 L of a 10% alcohol solution but only has a 30% alcohol solution. How many liters each of the 30 solution and water should be mixed to make the desired 18 liters of a 10% solution?
--------------
Equation:
water + water = water
0.70x + 1*(18-x) = 0.90*18
-----
70x + 100(18-x) = 90*18
-----
70x + 100*18 - 100x = 90*18
======
-30x = -10*18
x = 6 L (amt of 30% solution needed)
18-x = 18-6 = 12 L (amt of water needed)
----------------------
Cheers,
Stan H.
=============